gcc8617:

WHAT IS 18- 5/10+8/10

KynxKatz:

Mm one sec

KynxKatz:

If you want further explanation, kindly dm me

KynxKatz:

18−5/10+8/10 =18-1/2+8/10 =35/2+8/10 =35/2=4/5 =183/10

PWRDrEaMz:

Shouldn't there be parenthesis around any of the problem?